OpenStudy (anonymous):

can you help me solve- x/4+2x=1/3. Please...im sooo lost

OpenStudy (shamil98):

Cross multiplication works nicely with this.

OpenStudy (jdoe0001):

\(\bf \cfrac{x}{4}+2x=\cfrac{1}{3}\quad \textit{multiply both sides by 12}\\ \quad \\ 12\left(\cdot \cfrac{x}{4}+2x\right)=12\cdot \cfrac{1}{3}\) see what you get